OpenStudy (anonymous):

draw a graph of f(x). assume that f(x) is defined and continuous for all real x. Given, the Limit of f(x) ad x approaches infinity equals -infinity, and limit of f(x) as x approaches -infinity equals + infinity.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} = -\infty and \lim_{x \rightarrow -\infty}=+\infty\]